# Edge Delayed Deep Deterministic Policy Gradient: efficient continuous control for edge scenarios

### v1

@@ 5 Smoothing the optimization landscape
$L(\theta)=(r+\gamma Q_{\theta^{\prime}}(s^{\prime},\mu_{\phi^{\prime}}(s^{\prime}))-Q_{\theta}(s,a))^{2}+\xi||\nabla_{a}Q(s,a)||^{2}$ (14)
@@
[…]
@@
$\nabla_{\mu_{\phi}(s)}Q_{\theta}(s,\mu_{\phi}(s))\approx\nabla_{\mu_{\phi}(s)}Q^{*}(s,\mu_{\phi}(s))$. For this reason, we can take advantage of the delayed update introduced in [5]. This way, not only do we offer more time to the $Q_{\theta}$ to improve, but we also save computational resources, skipping $k-1$ updates every $k$, where $k$ is the frequency of actor updates.
@@
        $d=s^{\prime}\text{ is terminal}$
@@
        Sample mini-batch of $N$ tuples $(s,a,r,s^{\prime},d)$ from $\mathcal{B}$
@@
Figure 3: Training progress curves for continuous control tasks in OpenAI Gym, showing the effect of different choices of $\alpha,\beta$ in EdgeD3 compared to EdgeDDPG. Plots and shaded areas indicate mean and half a standard deviation, respectively, from evaluation across 10 trials. Benchmarks were performed on 10 random seeds for simulator and network initializations. Curves are smoothed uniformly for visual clarity.
6 Experiments
@@
For our implementation of all the algorithms, we used a feed-forward network composed of 3 layers of 256 neurons, optimized using Adam optimizer [25] with $3\cdot 10^{-4}$ stepsize for a fair comparison. For EdgeD3, TD3 and SAC we considered $k=2$, thus the actor is updated every $2$ updates of the Q-functions. For the target smoothing distribution, we use the proposed clipped Gaussian distribution also used in [5], and for the exploration policy, we used a Gaussian distribution $N(0,0.1)$ for all the algorithms, apart from SAC [6] where we used the learned posterior distribution. More technical details for reproducibility can be found in appendix A. In the rest of this section, we firstly compare the memory usage of the proposed algorithm with state of the art, secondly, their GPU-time utilization, and thirdly, we compare learning performance on the Mujoco benchmarks.
6.1 Resource use comparison
@@
6.1.1 RAM usage
The most popular algorithms that build upon DDPG, implementing the DPG estimator[23], are SAC and TD3, which both utilize the CDQ mechanism, exploiting an ensemble of two Q functions to estimate the Q-learning target. However, having this additional function also implies having another additional target function, thus effectively having four networks in total to maintain in memory. Such additional cost is justified by the improvements in performance. However, if memory is a concern, such as for edge computing, this additional cost might not be worth it. For this reason, we will compare the algorithms by their memory consumption. This section, in conjunction with section 6, shows we can achieve state-of-the-art performances with much less memory required, thanks to the new loss formulation.
Since all algorithms share the same Replay Buffer size, the only factor influencing the footprint is the number of networks that the algorithms require. In table 1, we show the percentage of decrease in peak memory usage of 10.000 update steps of each algorithm using 10-dimensional fake Gaussian noise data generated at the beginning, thus removing the replay buffer from the memory consumption, and also removing the environment, which might cause some sharp increases in memory usage biasing the estimates. The test was carried out using the CPU for the computation so that the memory used by the process was not split across RAM and GPU memory.
Table 1: Comparison of percentage of peak additional memory used compared to EdgeD3.
@@ 6.1.1 RAM usage
DDPG $-1.2\%$
TD3 $+29.3\%$
SAC $+31.1\%$
Figure 4: Comparing EdgeD3 with baselines in continuous control tasks. Plots are from 10 random seeds for simulator and network initializations, smoothed for visualization. Evaluations of Return are performed every 5000 time steps; plots show mean and half a standard deviation over 10 episodes.
6.1.2 GPU-time comparison
@@
Since all have a Replay Buffer $\mathcal{B}$, and that all at inference time have roughly the same cost, as the state is forwarded through the policy network $\mu_{\phi}$ and some Gaussian noise is added to the network prediction, the only component that can vary the computational cost, is the training loop. For this reason, we create random data from a $10$-dimensional Gaussian distribution and use that as fake states for the forward passes. We then proceed to run $10000$ training steps for algorithms and repeat it over $10$ different seeds. The results of the comparisons are reported in fig. 5. For clarity, we also report the improvements in table 2. Indeed, it’s evident how the proposed method, compared to the other methods, requires 30% less computing time since either they lack the delayed update or they require the update of 2 Q-functions.
Table 2: Comparison of GPU-time. Details on the hardware can be found in section 6.1. Between parenthesis is reported the percentage of time saved by EdgeD3 compared to the various methods.
@@
DDPG $285.5\pm 7.4$ms ($-25.0\%$)
TD3 $308.2\pm 2.7$ms ($-30.5\%$)
SACd $320.90\pm 3.6$ms ($-33.3\%$)
SACo $492.91\pm 2.9$ms ($-56.8\%$)
Figure 5: GPU time for one step of training loop for the different algorithms averaged over $10000$ steps, repeated over 10 different seeds. Details on the hardware can be found in section 6.1
@@
6.1.3 Comparison to state-of-the-art
For the comparison with other algorithms, as reported at the beginning of the section, we will use the Mujoco suit. For reproducibility, we used the same criterion used by the authors TD3 [5], having a bootstrap phase at the beginning of the learning. Each task ran for $1$ million steps and was evaluated every $5000$ step on 10 different environment initializations per evaluation. The reported results are also averaged over 10 different independent learning with a different seed each for different environment and network initializations.
@@ 6.1.3 Comparison to state-of-the-art
However, in order to have a fair comparison, instead of comparing the algorithms based on the environment steps, we compare them based on their training time. For this reason, since it’s the cheapest across the pool, we first run the EdgeD3 algorithm in all tasks, keeping track of the wall-clock time required to do so for each environment. Then, we proceed to run all the other algorithms for the same amount of time in order to take into consideration the additional computational cost brought by the double Q-function update shown in table 2.
@@
The final evaluations are reported in fig. 4. In table 3, we report the best score achieved by the various algorithms, allowing each the same amount of wall-clock time. In bold, we report the best algorithms, whereas the underlined ones are the two best.
@@
It can be seen how EdgeD3 is consistently part of the two best algorithms. Furthermore, it matches and even surpasses performances of state-of-the-art methods while having a significantly smaller memory footprint, as reported in table 1. Indeed, it can also be seen that is not perfect, as it still struggles on very complex tasks such as Humanoid. On the other hand, its memory-wise competitor, DDPG, is highly outperformed by EdgeD3 in almost all the tasks, even going from a non-convergence regime to a state-of-the-art policy.
@@
Table 3: Maximum average return over 10 evaluations across 10 trials, training a policy for the same amount of wall-clock time. The bolded values represent the best policy learned by each algorithm for each task, and the underlined ones represent the two best.
@@
7 Conclusions and future work
Edge computing is always gaining more attention thanks to its ability to allow for scalable deployment and preserving the privacy of the final user, handling the computation directly on-device.
We present Edge Delayed Deep Deterministic Policy Gradient (EdgeD3), which builds on top of Deep Deterministic Policy Gradient (DDPG) [4]. We introduce a new lightweight, easy-to-implement, highly tunable loss that trades off overestimation and underestimation by exploiting an unbalanced loss, described in eq. 11. Furthermore, we include new tricks to stabilize the training without additional costs. This algorithm aims at being a step towards scalable Deep Reinforcement Learning algorithms for edge computing, thus aiming at minimizing the computational cost and the memory footprint, while not hindering performances.
As done by previous works, such as TD3 and SAC, it achieves better performance by tackling the overestimation bias brought by the temporal difference loss of Q-learning. However, instead of using an ensemble of estimators, EdgeD3 exploits a new Expectile loss to do so while avoiding adding computational burden, thus keeping its property of being edge-friendly.
This new expectile loss, combined with additional blocks proposed by various other works in literature, such as the target smoothing and the delayed policy update, allow us to create a method that is $30\%$ computationally cheaper than the current state-of-the-art methods, preserves a memory footprint on the same level as the original algorithm and $30\%$ smaller that the state of the art, all by preserving the same performances of such more computationally demanding algorithms.
Potential future research avenues involve investigating other unbalanced losses, such as quantile loss or unbalanced-huber loss, which have the same computational cost as the proposed expectile loss while being more robust to outlier values. Furthermore, the proposed method introduces a hyperparameter that controls the overestimation and the underestimation. Even though a priori good guess for such hyperparameters exists, there is the possibility to extend the current algorithm with online fine-tuning of such hyperparameters in order to tune it automatically. Finally, we aim to test the presented method in real-world scenarios, thus leveraging its ability to be very computationally cheap and to have a small memory footprint to carry out learning directly on edge devices of several control tasks.
@@ Appendix A Reproducibility
A.1 Networks architectures
For TD3, DDPG, EdgeDDPG, and EdgeD3 the following network architecture have been used for the actor:
(state dim -> 256)
Relu
(256 -> 256)
Relu
(256 -> action dim)
tanh
For SAC, the following network architecture has been used for the actor:
(state dim -> 256)
Relu
(256 -> 256)
Relu
(256 -> action dim * 2)
For all methods, the following network architecture has been used for the critic:
(state dim + action dim-> 256)
Relu
(256 -> 256)
Relu
(256 -> 1)
A.2 Hyper-parameters

### v2

@@ V Smoothing the optimization landscape
[…]
@@
Indeed, $p(x)$ wants to be a smoothing function, having $y=\int p(\epsilon)[r+\gamma Q_{\theta^{\prime}}(s^{\prime},\mu_{\phi^{\prime}}+\epsilon(s^{\prime}))]d\epsilon$ so it’s suggested to choose a distribution $p$ centered in 0 and symmetric, in order to avoid introducing any bias.
@@
$\nabla_{\mu_{\phi}(s)}Q_{\theta}(s,\mu_{\phi}(s))\approx\nabla_{\mu_{\phi}(s)}Q^{*}(s,\mu_{\phi}(s))$. For this reason, we can take advantage of the delayed update introduced in [14]. This way, not only do we offer more time to the $Q_{\theta}$ to improve, but we also save computational resources, skipping k1k-1 updates every $k$, where $k$ is the frequency of actor updates.
@@
   $d=\begin{cases}1\text{ if }s^{\prime}\text{ is terminal}\\ 0\text{ otherwise}\end{cases}$
@@
   Sample batch of NN tuples $(s,a,r,s^{\prime},d)$ from $\mathcal{B}$
@@
Figure 3: Training progress curves for continuous control tasks in OpenAI Gym, showing the effect of different choices of $\alpha,\beta$ in EdgeD3 compared to EdgeDDPG. Plots and shaded areas indicate mean and standard deviation, respectively, from evaluation across 10 trials. Benchmarks were performed on 10 random seeds for simulator and network initializations. Curves are smoothed uniformly for visual clarity.
VI Experiments
@@
For our implementation of all the algorithms, we used a feed-forward network composed of 3 layers of 256 neurons, optimized using Adam optimizer [44] with $3\cdot 10^{-4}$ stepsize for a fair comparison. For EdgeD3, TD3 and SAC we considered $k=2$, thus the actor is updated every $2$ updates of the Q-functions. For the target smoothing distribution, we use the proposed clipped Gaussian distribution also used in [14], and for the exploration policy, we used a Gaussian distribution $N(0,0.1)$ for all the algorithms, apart from SAC [15] where we used the learned posterior distribution. More technical details for reproducibility can be found in appendix A. In the rest of this section, we firstly compare the memory usage of the proposed algorithm with state of the art, secondly, their GPU-time utilization, and thirdly, we compare learning performance on the Mujoco benchmarks. All results repoted come from networks trained using eq. 11 with $\alpha=1,\beta=2$, thus showing that very limited hyperparameter tuning is needed, and that a performant initial guess exists.
VI-A Resource use comparison
@@
VI-A1 RAM usage
The most popular algorithms that build upon DDPG, implementing the DPG estimator[42], are SAC and TD3, which both utilize the CDQ mechanism, exploiting an ensemble of two Q functions to estimate the Q-learning target. However, having this additional function also implies having another additional target function, thus effectively having four networks in total to maintain in memory. Such additional cost is justified by the improvements in performance. However, if memory is a concern, such as for edge computing, this additional cost might not be worth it. For this reason, we will compare the algorithms by their memory consumption. This section, in conjunction with section VI, shows we can achieve state-of-the-art performances with much less memory required, thanks to the new loss formulation.
Since all algorithms share the same Replay Buffer size, the only factor influencing the footprint is the number of networks that the algorithms require. In table I, we show the percentage of decrease in peak memory usage of 10.000 update steps of each algorithm using 10-dimensional fake Gaussian noise data generated at the beginning, thus removing the replay buffer from the memory consumption, and also removing the environment, which might cause some sharp increases in memory usage biasing the estimates. The test was carried out using the CPU for the computation so that the memory used by the process was not split across RAM and GPU memory.
TABLE I: Comparison of percentage of peak additional memory used compared to EdgeD3.
@@ VI-A1 RAM usage
DDPG $1.2\%$
TD3 $29.3\%$
SAC $31.1\%$
Figure 4: Comparing EdgeD3 with baselines in continuous control tasks. Plots are from 10 random seeds for simulator and network initializations, smoothed for visualization. Evaluations of Return are performed every 5000 time steps; plots show a standard deviation over 10 episodes.
VI-A2 GPU-time comparison
@@
Since all have a Replay Buffer $\mathcal{B}$, and that all at inference time have roughly the same cost, as the state is forwarded through the policy network $\mu_{\phi}$ and some Gaussian noise is added to the network prediction, the only component that can vary the computational cost, is the training loop. For this reason, we create random data from a $10$-dimensional Gaussian distribution and use that as fake states for the forward passes. We then proceed to run $10000$ training steps for algorithms and repeat it over $10$ different seeds. The results of the comparisons are reported in fig. 5. For clarity, we also report the improvements in table II. Indeed, it’s evident how the proposed method, compared to the other methods, requires 30% less computing time since either they lack the delayed update or they require the update of 2 Q-functions.
TABLE II: Comparison of GPU-time. Details on the hardware can be found in section VI-A. Between parenthesis is reported the percentage of time saved by EdgeD3 compared to the various methods.
@@
DDPG $285.5\pm 7.4$ms ($25.0\%$)
TD3 $308.2\pm 2.7$ms ($30.5\%$)
SACd $320.90\pm 3.6$ms ($33.3\%$)
SACo $492.91\pm 2.9$ms ($56.8\%$)
Figure 5: GPU time for one step of training loop for the different algorithms averaged over $10000$ steps, repeated over 10 different seeds. Details on the hardware can be found in section VI-A
@@
VI-B Comparison to state-of-the-art
For the comparison with other algorithms, as reported at the beginning of the section, we will use the Mujoco suit. For reproducibility, we used the same criterion used by the authors TD3 [14], having a bootstrap phase at the beginning of the learning. Each task ran for $1$ million steps and was evaluated every $5000$ step on 10 different environment initializations per evaluation. The reported results are also averaged over 10 different independent learning with a different seed each for different environments and network initializations.
@@ VI-B Comparison to state-of-the-art
However, in order to have a fair comparison, instead of comparing the algorithms based on the environment steps, we compare them based on their training time. For this reason, since it’s the cheapest across the pool, we first run the EdgeD3 algorithm in all tasks, keeping track of the wall-clock time required to do so for each environment. Then, we proceed to run all the other algorithms for the same amount of time in order to take into consideration the additional computational cost brought by the double Q-function update shown in table II.
@@
The final evaluations are reported in fig. 4. In table III, we report the best score achieved by the various algorithms, allowing each the same amount of wall-clock time. In bold, we report the best algorithms, whereas the underlined ones are the two best.
@@
It can be seen how EdgeD3 is consistently part of the two best algorithms. Furthermore, it matches and even surpasses performances of state-of-the-art methods while having a significantly smaller memory footprint, as reported in table I. Indeed, it can also be seen that is not perfect, as it still struggles on very complex tasks such as Humanoid. On the other hand, its memory-wise competitor, DDPG, is highly outperformed by EdgeD3 in almost all the tasks, even going from a non-convergence regime to a state-of-the-art policy.
VI-C Real world edge evaluation
To further validate the applicability of EdgeD3 beyond simulated environments, we conducted a set of real-world edge evaluations on a custom-built TurtleBot platform equipped with a 2D LiDAR sensor and an onboard Raspberry Pi3B device. The robot is a differential drive system, its control input $u_{t}$ is composed by the target linear and angular velocities, which are tracked by a low-level controller. This hardware configuration reflects a prototypical edge robotics scenario, where both compute and memory resources are constrained.
We evaluated the proposed algorithm on two robotic navigation tasks. For both settings, the state space, action space, and reward functions are shared. The state space consists of a downsampled vector of laser scan measurements combined with the linear and angular velocities of the robot:
$s_{t}=\begin{bmatrix}d_{1}&d_{2}&\ldots&d_{n}&v_{t}&\omega_{t}\end{bmatrix}^{T},$ (17)
where $d_{i}$ are evenly sampled laser range readings ($n=16$), $v_{t}$ and $\omega_{t}$ are the measured linear and angular velocities. The minimum laser reading $d_{\min}=\min\{d_{1},\ldots,d_{n}\}$ is used for collision detection and reward shaping.
The action space $a_{t}\in[-1,1]^{2}$ is used to control the robot following the control law:
$u_{t}$ $=B\bigl(M\,a_{t}+d\bigr)$ (18)
[…]
where $B\in\mathbb{R}^{2\times 2}$ is a diagonal matrix containing the maximum linear and angular velocities of the robot, $M\in\mathbb{R}^{2\times 2}$ is an affine-scaling matrix, and $d\in\mathbb{R}^{2}$ is a translation vector. This transformation constrains the robot to move only in the forward direction.
The reward function is designed to encourage fast, smooth motion while penalizing proximity to obstacles and excessive angular velocity:
$r_{t}$ $=\begin{cases}3v_{t}-\left|\frac{\omega_{t}}{2}\right|-\frac{1}{2}\left(1-d_{\min}\right),&\text{if }d_{\min}\geq 0.2\\ -5,&\text{otherwise}\end{cases}$ (21)
The episodes end when $d_{\text{min}}<0.2$ or a maximum time limit is reached. This reward encourages a reactive navigation that maximizes the travel distance, measured by integrating the linear velocity, while performing the least possible turns.
  1. 1.
    Corridor Navigation: In this task, the robot operates within a confined rectangular ring-shaped environment, fig. 6 (top). The episode time limit is $17$ seconds.
  2. 2.
    Unstructured navigation: In this task, the robot is deployed within a substantially larger and more complex indoor environment populated with a diverse array of static obstacles, shown in fig. 6 (bottom). The episode time limit is $25$ seconds.
At the beginning of the training sequences of both scenarios, the robot is manually placed in a random location. Then it runs autonomously until the maximum time is reached. When $d_{min}<0.2$, a custom routine repositions the agent using the complete laser scan for orientation, ensuring the robot is parallel to the closest vertical surface and $d_{min}>0.3$. No external localization system or odometry is employed in these experiments. In both tasks, the goal is to continuously move forward without collisions. The agent receives only real-time LiDAR and speed measurements as observations, and must learn a policy that maximizes the total traveled distance while maintaining safe navigation and avoiding obstacles.
Figure 6: First column shows real-world images of the two tasks, corridor navigation and unstructured navigation. The second column shows the training curves for 1 hour of training for all algorithms, averaged over 5 different seeds.
For both tasks, we compared EdgeD3 to the baselines DDPG, SAC, and TD3 policies, each implemented using the same three-layer, 64-neuron architecture, using a batch size of 128 samples, and all running under identical system constraints. All methods are given 1 wall-clock hour budget for both tasks. Each algorithm was executed for 4 different random seeds.
In both scenarios, shown in fig. 6, EdgeD3 obtains strong performance in learning, surpassing all other state-of-the-art methods. In particular, we observe similar relative update rates to the ones measured in the simulations (reported in table II). Therefore, given the same wall-clock time, EdgeD3 is able to perform more updates than all other approaches, thanks to its lightweight update loop.
For reference, on the robot’s hardware, the working frequency of the algorithms during training is, respectively $8Hz,5.9Hz,5.8Hz,3.3Hz$ for EdgeD3, TD3, DDPG, and SAC (without delayed updates). Recent algorithms discussed in Section II are not applicable to this setup.
A detailed examination of fig. 6 reveals distinct learning dynamics for the evaluated algorithms across the two real-world scenarios. In the corridor navigation task (Scenario 1), EdgeD3 demonstrates consistently superior performance from the outset. This is attributable to the relatively structured nature of the environment, where EdgeD3’s stabilized target estimation and efficient exploration rapidly yield an effective navigation policy with minimal initial exploration cost.
Conversely, in the unstructured navigation scenario (Scenario 2), EdgeD3 exhibits suboptimal performance during the initial training phase, with a marked transition to superior performance approximately 30 minutes into training. This phenomenon arises probably due to the increased environmental complexity and variability, which necessitate more extensive exploration and model adaptation before effective generalization can occur. During the initial phase, EdgeD3’s conservative target updates, while stabilizing, may delay the exploitation of promising policies as the algorithm accumulates sufficient state-action coverage.
@@ VI-C Real world edge evaluation
TABLE III: Maximum average return over 10 evaluations across 10 trials, training a policy for the same amount of wall-clock time. The bolded values represent the best policy learned by each algorithm for each task, and the underlined ones represent the two best.
@@
VII Conclusions and future work
Edge computing is always gaining more attention thanks to its ability to allow for scalable deployment and preserving the privacy of the final user, handling the computation directly on-device. We present Edge Delayed Deep Deterministic Policy Gradient (EdgeD3), which builds on top of Deep Deterministic Policy Gradient (DDPG) [13]. We introduce a new lightweight, easy-to-implement, highly tunable loss that trades off overestimation and underestimation by exploiting an unbalanced loss, described in eq. 11. Furthermore, we include new tricks to stabilize the training without additional costs. This algorithm aims to be a step towards scalable Deep Reinforcement Learning algorithms for edge computing, thus aiming at minimizing the computational cost and the memory footprint, while not hindering performance. As done by previous works, such as TD3 and SAC, it achieves better performance by tackling the overestimation bias brought by the temporal difference loss of Q-learning. However, instead of using an ensemble of estimators, EdgeD3 exploits a new Expectile loss to do so while avoiding adding computational burden, thus keeping its property of being edge-friendly.
This new expectile loss, combined with additional blocks proposed by various other works in literature, such as the target smoothing and the delayed policy update, allow us to create a method that is $30\%$ computationally cheaper than the current state-of-the-art methods, preserves a memory footprint on the same level as the original algorithm and $30\%$ smaller than the state of the art, all by preserving the same performances of such more computationally demanding algorithms. Furthermore, the approach is evaluated on real-world edge scenarios, validating its ability to learn effective policies with minimal computational resources.
Potential future research avenues involve investigating other unbalanced losses, such as quantile loss or unbalanced-huber loss, which have the same computational cost as the proposed expectile loss while being more robust to outlier values. Furthermore, the proposed method introduces a hyperparameter that controls the overestimation and the underestimation. Even though an a priori good guess for such hyperparameters exists, there is the possibility to extend the current algorithm with online fine-tuning of such hyperparameters in order to tune it automatically.
@@ Appendix A Reproducibility
A-A Practical Considerations for the Selection of $\alpha$ and $\beta$
The hyperparameters $\alpha$ and $\beta$ play pivotal roles in the stability and efficacy of the proposed algorithm. In practice, $\alpha$ and $\beta$ are used to balance the contributions of two competing estimators within the target update mechanism. The regime $\alpha<\beta$ is generally preferred in environments characterized by high stochasticity or when the policy exhibits high variance, as it introduces additional conservatism in target value estimation, thereby mitigating the risk of overestimation bias. This setting is often advantageous in real-world robotics and control scenarios where measurement noise or model uncertainty is prevalent.
Conversely, the regime $\alpha>\beta$ can be considered in highly deterministic environments or when empirical evidence suggests persistent underestimation of value functions, potentially leading to overly conservative policies. In these settings, a slight relaxation of conservatism may facilitate more aggressive exploration and policy improvement. However, in our experience, such cases are infrequent, and $\alpha\leq\beta$ remains the default configuration for most applications. This regime can be thought of as a loss-induced optimism in the face of uncertainty, famous in RL to improve exploration.
Both $\alpha$ and $\beta$ are treated as tunable hyperparameters and should be selected via standard validation procedures, e.g., grid search or Bayesian optimization, within a plausible range. It is crucial to monitor both the learning stability and the empirical performance during tuning. We emphasize that the primary determinant for $\alpha$ and $\beta$ selection should be the desired trade-off between optimism and conservatism in target value estimation, guided by the stochasticity and complexity of the deployment environment. Default values of $\alpha=1$ and $\beta=2$ have demonstrated robust performance across a wide variety of tasks in our evaluations, therefore are highly suggested as a starting point for the tuning, if even needed.
A-B Hyper-parameters and further ablations